A 250 , 25-cm-diameter plastic disk is spun on anaxle through its center by an electric motor. What torque must th...

A 250 g, 25-cm-diameter plastic disk is spun on anaxle through its center by an electric motor.
What torque must the motor supply to take the disk from 0to 1600 rpm in 4.8 s?
Express your answer using two significant figures.(answer?=N*m)
Answer #1
Concepts and reason

The concepts used to solve this problem are angular kinematics and moment of inertia.

Initially, the angular acceleration can be calculated by using the angular kinematics formula. Then, the moment of inertia can be calculated for the disk by using its formula. Finally, the torque acting on the system can be calculated by multiplying the moment of inertia and angular acceleration.

Fundamentals

The expression for the angular acceleration from angular kinematics is,

ωf=ωi+αt{\omega _f} = {\omega _i} + \alpha t

Here, ωf{\omega _f} is the final angular speed, ωi{\omega _i} is the initial angular speed, α\alpha is the angular acceleration, and tt is the time period.

The moment of inertia for the disk is,

I=12MR2I = \frac{1}{2}M{R^2}

Here, II is the moment of inertia, MM is the mass of the disk, and RR is the radius of the disk.

The formula for the torque is,

τ=Iα\tau = I\alpha

Here, τ\tau is the torque acting on the system.

Rpm is revolution per min which is expressed as rpm=revmin{\rm{rpm}} = \frac{{{\rm{rev}}}}{{\min }} .

The expression for the angular acceleration is,

ωf=ωi+αt{\omega _f} = {\omega _i} + \alpha t

Substitute 1600rpm1600\;{\rm{rpm}} for ωf{\omega _f} , 00 for ωi{\omega _i} , 4.8s for tt .

0+4.8α=(1600revmin)(1min60s)(2πrad)167.47=4.8αα=34.9rad/s2\begin{array}{c}\\0 + 4.8\alpha = \left( {1600\,\frac{{{\rm{rev}}}}{{\min }}} \right)\left( {\frac{{1\min }}{{60\,{\rm{s}}}}} \right)\left( {2\pi \,{\rm{rad}}} \right)\\\\167.47 = 4.8\alpha \\\\\alpha = 34.9\,{\rm{rad/}}{{\rm{s}}^2}\\\end{array}

The radius of the plastics disk is,

R=d2R = \frac{d}{2}

Here, d is the diameter of the disk.

Substitute 25cm25\,{\rm{cm}} for d to find R.

R=(25cm)(102m1cm)2=0.125m\begin{array}{c}\\R = \frac{{\left( {25\,{\rm{cm}}} \right)\left( {\frac{{{{10}^{ - 2}}\,{\rm{m}}}}{{1\,{\rm{cm}}}}} \right)}}{2}\\\\ = 0.125\,{\rm{m}}\\\end{array}

The expression for the moment of inertia is,

I=12MR2I = \frac{1}{2}M{R^2}

Substitute 250g250{\rm{g}} for MM and 0.125m0.125\,{\rm{m}} for RR .

I=12(250g)(1×103kg1g)[(12.5cm)(1×102m1cm)]2=1.95×103kgm2\begin{array}{c}\\I = \frac{1}{2}\left( {250\,{\rm{g}}} \right)\left( {\frac{{1 \times {{10}^{ - 3}}\,{\rm{kg}}}}{{1\,{\rm{g}}}}} \right){\left[ {\left( {12.5\,{\rm{cm}}} \right)\left( {\frac{{1 \times {{10}^{ - 2}}\,{\rm{m}}}}{{1\,{\rm{cm}}}}} \right)} \right]^2}\\\\ = 1.95 \times {10^{ - 3}}\;{\rm{kg}} \cdot {{\rm{m}}^{\rm{2}}}\\\end{array}

The expression for the torque is,

τ=Iα\tau = I\alpha

Substitute 1.95×103kgm21.95 \times {10^{ - 3}}\,{\rm{kg}} \cdot {{\rm{m}}^{\rm{2}}} for II and 34.9rad/s234.9\,{\rm{ rad/}}{{\rm{s}}^2} for α\alpha .

τ=(1.95×103kgm2)(34.9rad/s2)=0.0681Nm0.068Nm\begin{array}{c}\\\tau = \left( {1.95 \times {{10}^{ - 3}}\,{\rm{kg}} \cdot {{\rm{m}}^2}} \right)\left( {34.9\,{\rm{rad/}}{{\rm{s}}^2}} \right)\\\\ = 0.0681\,{\rm{N}} \cdot {\rm{m}}\\\\ \approx {\rm{0}}{\rm{.068}}\,{\rm{N}} \cdot {\rm{m}}\\\end{array}

Ans:

The torque acting on the system is 0.068Nm0.068\;{\rm{N}} \cdot {\rm{m}} .